OpenStudy (anonymous):

is 72 multiple of of 10 ? explain

OpenStudy (turingtest):

you mean a whole number multiple?

OpenStudy (turingtest):

make a list: 10*1=10 10*2=20 10*3=30 etc. if you find that the list skips 72, then the answer is no if 72 show up on the list the answer is yes